OpenStudy (anonymous):

what are the sine, cosine, and tangent of theta=3pi/4 radians

OpenStudy (jdoe0001):

use the Unit Circle :)

OpenStudy (anonymous):

can you help me use it? it confuses me :/

OpenStudy (jdoe0001):

well, what's the "x" and "y" values for \(\bf \cfrac{3\pi}{4}\quad?\)

OpenStudy (anonymous):

im not sure, how can i figure that out?

OpenStudy (anonymous):

i

OpenStudy (anonymous):

is x the 3pi?

OpenStudy (jdoe0001):

well, no quite, \(\bf \cfrac{3\pi}{4}\) is the angle

OpenStudy (anonymous):

@jdoe0001 please help me real fast :/

OpenStudy (anonymous):

oh good sorry my screen glitched xD made it look like you logged off

OpenStudy (anonymous):

the angle is 135 right?

OpenStudy (jdoe0001):

hmm use -> http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_lg.gif <-- there's a good unit circle there then just use the sine, cosine and tangent identities, notice the pair there,are the x, y pair keep in mind that \(\bf sin(\theta)=\cfrac{y}{hypotenuse}\qquad cos(\theta)=\cfrac{x}{hypotenuse}\qquad tan(\theta)\cfrac{y}{x}\)

OpenStudy (jdoe0001):

yes \(\bf \bf \cfrac{3\pi}{4} \implies 135^o\)

OpenStudy (jdoe0001):

and of course, the hypotenuse for a Unit circle, is 1, or the radius

OpenStudy (anonymous):

D. sin theta=- square root 2/2; cos theta =square root 2/2; tan theta =-1. is that right? i tried using your unit circle to get it

OpenStudy (anonymous):

wait would tan be 1?

OpenStudy (jdoe0001):

nope , is not tangent is not posiitive, 2nd quadrant has negative "x" and positive "y" thus y/x will yield a negative so sin theta=- square root 2/2; cos theta =square root 2/2; tan theta =-1. \(\Large \checkmark\)

OpenStudy (anonymous):

oh so the answer would be D?

OpenStudy (anonymous):

Thanks for the help :D

OpenStudy (jdoe0001):

if that's what D has, then yes,

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i got it wrong :/ i put what you said too